Circular No. 8496

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005at AND 2005au Further to IAUC 8490, R. Martin reports the discovery of an apparent

supernova (mag 16.0) on CCD images taken on Mar. 15.83 and 16.67 UT. The new object is located at α = 19^{h}09^{m}53^{s}.62, δ = −63^{o}49′24′′.1 (equinox 2000.0). Nothing was visible at this location in an image from 2004 Aug. 8.75 (limiting mag 18.5). H. Yamaoka, Kyushu University, adds that there is a foreground star (mag ∼ 18) located ∼ 5′′ south of Martin’s position. L. A. G. Monard, Pretoria, S. Africa, reports his independent discovery of SN 2005at on CCD images taken on Mar. 5.139 (at mag 14.3) and 19.133 (mag 14.4), providing position end figures 53^{s}.57, 22′′.8 (or 46′′ east and 117′′ north of the center of NGC 6744). Monard adds that nothing is visible at this location on his image from 2004 Feb. 15.1 (limiting red mag 17.3) or on an image from the Digitized Sky Survey (limiting red mag 20).

B. Schmidt and M. Salvo, Australian National University (ANU), report that a spectrogram (range 320–1050 nm) of SN 2005at, obtained on Mar. 19.77 UT with the ANU 2.3-m telescope (+ DBS), reveals it to be a type-Ic supernova resembling SN 1994I at ≈ 2 weeks past maximum light. The relatively red spectrum is dominated by broad P-Cyg profiles of the Ca II infrared triplet, Na D line, and what appears to be He I 1083-nm.

R. Arbour, South Wonston, U.K., reports his discovery of an apparent supernova (mag ≈ 15.8) on unfiltered CCD images taken with a 0.3-m reflector on Mar. 19.955 and 20.817 UT in the course of his supernova patrol. SN 2005au is located at α = 13^{h}16^{m}12^{s}.42, δ = +30^{o}56′40′′.5 (equinox 2000.0), which is 1′′ east and 21′′ south of the center of NGC 5056. Nothing is visible at this location on Digitized Sky Survey images from 1989 (limiting red mag 20.9) and 1995 (limiting blue mag 20.8). Arbour adds that images taken by D. Briggs (Clanfield, U.K., 0.60-m reflector) on Mar. 20.891 yield position end figures 12^{s}.38, 42′′.3 for SN 2005au.

CATACLYSMIC VARIABLE IN PYXIS = POSSIBLE NOVA IN PYXIS

T. Kawabata, Y. Kawabata, and K. Ayani, Bisei Astronomical Observatory (BAO); and H. Yamaoka, Kyushu University, obtained a lowdispersion (resolution 1000) spectrogram of the eruptive variable in Pyxis (cf. IAUC 8495) with the BAO 1.01-m telescope on Mar. 18.61 UT. The spectrum shows a blue continuum with wide absorption of Hβ and Hγ. The Hα feature is not clear, probably because the emission fills up the absorption. Other features are not clearly seen. This suggests that the object is not a classical nova, but rather a dwarf nova in outburst.

2005 March 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green